A computer company uses the linear model y = -21x + 13,082 to predict the decline in sales of a computer after it has been on the market more than one year. If x is the number of months after the first year and y is the number of computers sold during that month, how many computers will be sold 11 months after the first year?
1
Expert's answer
2019-04-18T09:21:04-0400

The number of computers sold each month after the first year is a decreasing arithmetic progression.


The first term of the progression (sales in first month after the first year) is:



"m_{1}=13082-21*1=13061"


The general view of the progression term is:



"m_{n}=13061-21*(n-1)"


The sum of sales for n is:


"S_{n}=\frac {2*m_{1}-(n-1)*21} {2}*n"


The sum of sales for 11 monthes is:


"S_{11}=\frac {2*13061-(11-1)*21} {2}*11=142516"




Answer: 142516 computers.